# Semiclassical descriptions of dissipative dynamics of strongly interacting Bose gases in optical lattices

### v2

@@ IV.3 Restored superfluid coherence after turning off the PA light
Figure 5: (a) SU(3) dTWA results for the non-unitary time evolution of the zero momentum distribution $S({\bf k}={\bf 0},t)$, where $\gamma$ varies from $\gamma\neq 0$ (indicated in the figure) to $\gamma=0$ at $t=\tau$, mimicking the abrupt turning off of the PA light at $t=\tau$. For $t\leq\tau$, the linear sweep time sequence is applied, as in Fig. 4, while the hopping $J(t)=J_{\rm fin}$ for $t>\tau$. The vertical dotted line indicates the endpoint of the sweep at $t=\tau$. (b) Same as (a) but for the SU(3) FP-TWA results.
@@
Figure 6: The restored peak height $\delta h=L^{-3}(S_{\rm peak}-S_{\rm fin})$ after turning off the PA laser, i.e., by setting $\gamma=0$ at the endpoint of the sweep (see Fig. 5), vs. the dissipation strength $\gamma$. The results are evaluated by (a) the SU(3) dTWA and (b) the SU(3) FP-TWA.
@@
Finally, we examine whether the SU(3) TWA can capture the restoration of the superfluid coherence after suddenly turning off the PA laser at $t=\tau$. A similar setup was implemented for the lattice ramp-down experiment in Ref. Tomita et al. (2017). It was observed in the experiment that after a certain hold time of the gas, an interference pattern in TOF images is restored, implying the phase coherence reformation Tomita et al. (2017). This suggests that the suppression of superfluid coherence is not attributed to the heating caused by the PA laser and that the final state after the ramping down remains a Mott insulating state.
@@
Figure 5 shows the SU(3) dTWA and FP-TWA results simulating the time evolution when $\gamma$ abruptly changes from $\gamma\neq 0$ to zero at $t=\tau$. We find that the increase of the zero momentum distribution occurs for all $\gamma$ studied after turning off the dissipation, implying the restoration of the phase coherence over long distances, as observed experimentally in the visibility measurements of TOF images Tomita et al. (2017). In the time region where $\gamma=0$, i.e., $t>\tau$, the momentum distribution at ${\bf k}={\bf 0}$ first increases up to a maximum peak, and then eventually saturates. Within the SU(3) TWA, the long time evolution for $t\gg\tau$ is expected to lead to a classical thermal state described by the Gibbs distribution for the classical chaotic Hamiltonian Wurtz et al. (2018), i.e., the SU(3) representation of the Bose-Hubbard model with a conserved filling factor.
@@
In order to better characterize the properties of the restored peaks in the momentum distribution, let us define $S_{\rm peak}=\max_{t>\tau}[S({\bf k}={\bf 0},t)]$ and $S_{\rm fin}=S({\bf k}={\bf 0},t=\tau)$. Figure 6 displays the restored peak height $\delta h=L^{-3}(S_{\rm peak}-S_{\rm fin})$ as a function of $\gamma$. In the SU(3) dTWA results shown in Fig. 6(a), the peak height $\delta h$ increases with increasing $\gamma$ for $\gamma\gtrsim 2$, while it decreases for $\gamma\lesssim 2$. The increase of $\delta h$ for $\gamma\gtrsim 2$ implies that many more particles stay in the system due to the stronger dissipation, i.e., the Zeno effect, before switching off the dissipation $\gamma$. Given the results in Figs. 4(a) and 4(b) on the atomic density in the linear sweep time sequence, the restored peak height $\delta h$ turns out to follow the total number of particles surviving in the optical lattice at the end point of the sweep, i.e., $t=\tau$.
@@
As shown in Fig. 6(b), a similar behavior of the $\gamma$ dependence of the peak height is also found in the SU(3) FP-TWA. Namely, $\delta h$ first decreases with increasing $\gamma$ and then starts to increase at a minimum point located around $1\leq\gamma\leq 2$. The presence of the minimum in this regime is also correlated to the behavior of the final state atom densities at $t=\tau$ shown in Figs. 4(c) and 4(d). However, we should note that the final states at $t=\tau$ are not reliable, even qualitatively, in the SU(3) FP-TWA, as implied by the numerically exact results in Sec. IV.4, and therefore we consult the SU(3) dTWA results to be compared with the experiment.
@@ IV.4 Numerically exact results
When the hopping amplitude is zero, i.e., $J=0$, the SU(3) dTWA reproduces the exact quantum time evolution generated by the master equation with a direct product state for the typical initial state. This is attributed to the linearity of the classical equation of motion in the interaction $U$ term and the dissipation $\gamma$ term as implied in Eq. (10). Figure 7(a) displays numerical results calculated by the SU(3) dTWA for a single site system, demonstrating that this specific semiclassical representation successfully reproduces the exact solutions of the dynamics of $\langle{\hat{X}}_{3}(t)\rangle={\rm Tr}[\rho_{\rm eff}(t)T_{3}]$ for arbitrary dissipation strength at all time.
However, as shown in Fig. 7(b), the quantitative validity of the SU(3) FP-TWA is verified only within a finite timescale. This limitation is due to the nonlinearity in the dissipation term of the Langevin equation. For $\gamma=0.5$, the SU(3) FP-TWA result coincides well with the numerically exact solution until $tU/\hbar\sim 3$. After this timescale, the SU(3) FP-TWA result deviates from the numerically exact solution and finally converges to a false steady state with $\langle{\hat{X}}_{3}(t)\rangle<1$ for $tU/\hbar\gg 1$. Note that the exact steady state for the present setup is the trivial bosonic vacuum state $|0\rangle$, implying $\langle{\hat{X}}_{3}(t)\rangle=1$ for $tU/\hbar\gg 1$. For larger values of $\gamma$, i.e., $\gamma=1$, $2$, and $4$, the threshold time is estimated as $tU/\hbar\sim 1.5$, $1$, and $0.5$, respectively. Hence, the valid timescale of this approximation shortens with increasing $\gamma$. Notice also that the time-evolved states for these different values of $\gamma$ result in the same false steady state at long time. Comparing the results in Figs. 7(a) and  7(b), the local linearity of the dissipative force in the SU(3) dTWA is essential to reproduce the correct steady state reached via the successive onsite losses over time.
@@
Figure 8: The SU(3) dTWA (left panels) and the FP-TWA (right panels) results of the non-unitary time evolution of the atomic density $n_{j}(t)$ for the dissipative Bose-Hubbard model on a four site ring (solid lines). The initial state at $t=0$ for the time evolution is the unit filling Mott insulating state $|\psi_{\rm Mott}\rangle$. The hopping amplitude $J$ is the same as in Fig. 2, i.e., $J=0.08U$, and the values of the dissipation strength $\gamma$ are indicated in the figures. For comparison, the numerically exact solutions are also plotted by dashed lines. In the SU(3) dTWA, the time grid size for the explicit fourth-order Runge-Kutta scheme is set to $\Delta t=2.0\times 10^{-2}\hbar/U$ and the number of generated classical trajectories is 20 0002000020\;00020 000. In the SU(3) FP-TWA, the grid size for the first-order Euler-Maruyama scheme is set to $\Delta t=2.0\times 10^{-4}\hbar/U$ and the number of generated Langevin trajectories is 200 000200000200\;000200 000.
Figure 9: Two body loss rate $\kappa$ as a function of $\gamma$ extracted from the non-unitary dynamics of the four system in Fig. 8. The time unit for the loss rate is $\hbar/U$ and the data points of $n_{j}(t)$ used for fitting in the form $(1+\kappa t)^{-1}$ are separated by $5\hbar/U$. The exact (red) and SU(3) dTWA (blue) results show a maximum at $\gamma\sim 2$, while the SU(3) FP-TWA results (orange) exhibit a maximum at $\gamma\sim 1$ within the standard fitting errors. To capture the early time behavior, the fitting is performed for the corresponding data $n_{j}(t)$ satisfying $t\leq 40\hbar/U$.
Figure 10: (a) Same as Figs. 4(a) and 4(b) but for a four-site ring. The number of generated classical trajectories is 20 0002000020\;00020 000. (b) Same as (a) but for the numerically exact solutions. (c) Same as (a) but for the SU(3) FP-TWA results. The number of the generated Langevin trajectories is 200 000200000200\;000200 000.
Reflecting these consequences, we also find in Fig. 9 that the two body loss rate $\kappa$ extracted from the dTWA results exhibits a maximum around at $\gamma\sim 2$, although the values themselves are smaller than the exact ones. However, the agreement with the three dimensional experiment in Fig. 2(c) suggests that the deviations between these two results of the loss rate may be smaller in higher dimensions, because the valid timescale of the dTWA increases with the connectivity between sites, i.e., the coordination number $z=2d$ Kunimi et al. (2021).
The right panels of Fig. 8 show the SU(3) FP-TWA results of the atomic density $n_{j}(t)$ compared with the numerically exact solutions. For the weak dissipation strength, i.e., $\gamma=0.1$, the SU(3) FP-TWA result deviates form the exact solution around $tU/\hbar=2$. However, for the strong dissipation strength such as $\gamma=4$ and $5$, the SU(3) FP-TWA and exact results coincide rather well until $tU/\hbar\lesssim 6$. We should note that this coincidence is not controlled in this approximation. Although the SU(3) FP-TWA can capture the continuous Zeno effect for the strong dissipation strength, it estimates the threshold at $\gamma\sim 1$, which is smaller than the exact solution. As shown in Fig. 9, the maximum of the two body loss rate $\kappa$ appears around the same value of $\gamma$. Comparing the SU(3) dTWA and FP-TWA results, we conclude that the qualitative performance of the SU(3) FP-TWA is worse than that of the SU(3) dTWA at long time.
@@ V Conclusions
In conclusions, we have generalized the SU(3) TWA method for the Bose-Hubbard system to include the Lindblad dynamics in the presence of onsite dissipation terms. We introduced two semiclassical representations of the dynamics governed by the effective Lindblad master equation: the discrete TWA representation and the Fokker-Planck TWA representation. These methods were applied to the dissipative Bose-Hubbard system with the two-body losses, previously studied experimentally in Ref. Tomita et al. (2017). We demonstrated that these semiclassical representations can qualitatively capture the essential features of the continuous quantum Zeno effect in the dissipative non-unitary dynamics, as observed experimentally. In particular, the discrete TWA representation was found to produce reasonable thresholds for the dissipation strength, distinguishing between the Zeno and non-Zeno regimes, which are comparable to the experimental results of the analog quantum simulation Tomita et al. (2017). This is attributed to the onsite linearity of the classical equation of motion in the discrete TWA representation, which was used in the Monte Carlo simulations.
As a future study, it would be interesting to apply these SU(3) TWA methods to other experiments associated with dissipative Bose-Hubbard systems Labouvie et al. (2016); Bouganne et al. (2020) and discuss the reproducibility of nontrivial driven-dissipative dynamical phenomena reported therein. The generalization of the methods to SU(${\cal N}$) spin systems is straightforward, and these methods potentially enable one to evaluate the impact of dissipation on various experiments of analog quantum simulation, such as optical tweezer arrays loaded with Rydberg gases, trapped ions, and large-$S$ dipolar gases in optical lattices. Beyond the present capability of the methods based on onsite variables, it is also a promising direction to incorporate the ideas of the cluster TWA with fluctuating cluster variables Wurtz et al. (2018) to treat offsite quantum correlations caused by the hopping term more accurately Nagao and Yunoki (2024).
@@ Acknowledgements.
A part of numerical simulations was carried out on the HOKUSAI supercomputing system at RIKEN (Project ID Q22576) and the supercomputer Fugaku installed at the RIKEN Center for Computational Science. We thank Kazuhiro Seki for his helpful comments on MPI parallelization for Monte Carlo simulations. We also thank Masaya Kunimi and Tetsuro Nikuni for fruitful discussions. We acknowledge Takafumi Tomita and Yoshiro Takahashi for sharing the experimental data in Ref. Tomita et al. (2017). This work was financially supported by JSPS KAKENHI (Grant Nos. JP18H05228, JP21H01014, JP21H04446, and JP21H03455), by MEXT Q-LEAP (Grant No. JPMXS0118069021), by JST FOREST (Grant No. JPMJFR202T), and by JST COI-NEXT (Grant No. JPMJPF2221).

### v3

@@ IV.3 Restored superfluid coherence after turning off the PA light
Figure 6: (a) SU(3) dTWA results for the nonunitary time evolution of the zero momentum distribution $S({\bf k}={\bf 0},t)$, where $\gamma$ varies from $\gamma\neq 0$ (indicated in the figure) to $\gamma=0$ at $t=\tau$, mimicking the abrupt turning off of the PA light at $t=\tau$. For $t\leq\tau$, the linear sweep time sequence is applied, as in Fig. 5, while the hopping $J(t)=J_{\rm fin}$ for $t>\tau$. The vertical dotted line indicates the endpoint of the sweep at $t=\tau$. (b) Same as (a) but for the SU(3) FP-TWA results.
@@
Figure 7: The restored peak height $\delta h=L^{-3}(S_{\rm peak}-S_{\rm fin})$ after turning off the PA laser, i.e., by setting $\gamma=0$ at the end point of the sweep (see Fig. 6), vs the dissipation strength $\gamma$. The results are evaluated by (a) the SU(3) dTWA and (b) the SU(3) FP-TWA.
@@
Finally, we examine whether the SU(3) TWA can capture the restoration of the superfluid coherence after suddenly turning off the PA laser at $t=\tau$. A similar setup was implemented for the lattice ramp-down experiment in Ref. Tomita et al. (2017). It was observed in the experiment that after a certain hold time of the gas, an interference pattern in TOF images is restored, implying the phase coherence reformation Tomita et al. (2017). This suggests that the suppression of superfluid coherence is not attributed to the heating caused by the PA laser and that the final state after the ramping down remains a Mott-insulating state.
@@
Figure 6 shows the SU(3) dTWA and FP-TWA results simulating the time evolution when $\gamma$ abruptly changes from $\gamma\neq 0$ to zero at $t=\tau$. We find that the increase of the zero-momentum distribution occurs for all $\gamma$ studied after turning off the dissipation, implying the restoration of the phase coherence over long distances, as observed experimentally in the visibility measurements of TOF images Tomita et al. (2017). In the time region where $\gamma=0$, i.e., $t>\tau$, the momentum distribution at ${\bf k}={\bf 0}$ first increases up to a maximum peak, and then eventually saturates. Within the SU(3) TWA, the long-time evolution for $t\gg\tau$ is expected to lead to a classical thermal state described by the Gibbs distribution for the classical chaotic Hamiltonian Wurtz et al. (2018), i.e., the SU(3) representation of the Bose-Hubbard model with a conserved filling factor.
@@
In order to better characterize the properties of the restored peaks in the momentum distribution, let us define $S_{\rm peak}=\max_{t>\tau}[S({\bf k}={\bf 0},t)]$ and $S_{\rm fin}=S({\bf k}={\bf 0},t=\tau)$. Figure 7 displays the restored peak height $\delta h=L^{-3}(S_{\rm peak}-S_{\rm fin})$ as a function of $\gamma$. In the SU(3) dTWA results shown in Fig. 7(a), the peak height $\delta h$ increases with increasing $\gamma$ for $\gamma\gtrsim 2$, while it decreases for $\gamma\lesssim 2$. The increase of $\delta h$ for $\gamma\gtrsim 2$ implies that many more particles stay in the system due to the stronger dissipation, i.e., the Zeno effect, before switching off the dissipation $\gamma$. Given the results in Figs. 5(a) and 5(b) on the atomic density in the linear sweep time sequence, the restored peak height $\delta h$ turns out to follow the total number of particles surviving in the optical lattice at the end point of the sweep, i.e., $t=\tau$.
@@
As shown in Fig. 7(b), a similar behavior of the $\gamma$ dependence of the peak height is also found in the SU(3) FP-TWA. Namely, $\delta h$ first decreases with increasing $\gamma$ and then starts to increase at a minimum point located around $1\leq\gamma\leq 2$. The presence of the minimum in this regime is also correlated to the behavior of the final state atom densities at $t=\tau$ shown in Figs. 5(c) and 5(d). However, we should note that the final states at $t=\tau$ are not reliable, even qualitatively, in the SU(3) FP-TWA, as implied by the numerically exact results in Sec. IV.4, and therefore we consult the SU(3) dTWA results to be compared with the experiment.
@@ IV.4 Numerically exact results
Figure 9: SU(3) dTWA (left panels) and FP-TWA (right panels) results of the nonunitary time evolution of the atomic density $n_{j}(t)$ for the dissipative Bose-Hubbard model on a four-site ring (solid lines). The initial state at $t=0$ for the time evolution is the unit filling Mott-insulating state $|\psi_{\rm Mott}\rangle$. The hopping amplitude $J$ is the same as in Fig. 2, i.e., $J=0.08U$, and the values of the dissipation strength $\gamma$ are indicated in the figures. For comparison, the numerically exact and mean-field results are also plotted by dashed and dotted lines, respectively. In the SU(3) dTWA, the time grid size for the explicit fourth-order Runge-Kutta scheme is set to $\Delta t=2.0\times 10^{-2}\hbar/U$ and the number of generated classical trajectories is 20 0002000020\;00020 000. In the SU(3) FP-TWA, the grid size for the first-order Euler-Maruyama scheme is set to $\Delta t=2.0\times 10^{-4}\hbar/U$ and the number of generated Langevin trajectories is 200 000200000200\;000200 000.
Figure 10: (a) Two-body loss rate $\kappa$ as a function of $\gamma$ extracted from the nonunitary dynamics of the four-site system in Fig. 9. The time unit for the loss rate is $\hbar/U$ and the data points of $n_{j}(t)$ used for fitting in the form $f_{\kappa}(t)=(1+\kappa t)^{-1}$ are separated by $5\hbar/U$. The exact (red) and SU(3) dTWA (blue) results show a maximum at $\gamma\sim 2$, while the SU(3) FP-TWA results (orange) exhibit a maximum at $\gamma\sim 1$ within the standard fitting errors. The mean-field result (black) also shows a maximum at $\gamma=2$. To capture the early-time behavior, the fitting is performed on the corresponding data $n_{j}(t)$ that satisfy $t\leq 40\hbar/U$. (b) Power-law exponents for the long-time behavior of the atomic density $n_{j}(t)$ for varying values of $\gamma$. The fitting range of time $t$ using the function $f_{\rm power}(t)=b(tU/\hbar)^{-c}$ is restricted to $tU/\hbar>40$. The values of $b$ and $c$ are rescaled by their corresponding maximum values, $b_{\rm max}$ and $c_{\rm max}$, where $(b_{\rm max},c_{\rm max})=(39.6,1.3)$, $(8,0.8)$, and $(3.1,0.5)$ for the exact, SU(3) dTWA, and SU(3) FP-TWA results, respectively. (c) Power-law fitting using the function $f_{\rm power}(t)$ and short-time fitting using the function $f_{\kappa}(t)$ for the atomic density $n_{j}(t)$ obtained by the numerically exact calculations for $\gamma=2$. For comparison, the mean-field result is also shown by a dashed line. (d) Same as (c), but for the results obtained from the dTWA. (e) Same as (c), but for the results obtained from the FP-TWA.
Figure 11: (a) Same as Figs. 5(a) and 5(b) but for a four-site ring. The number of generated classical trajectories is 20 0002000020\;00020 000. (b) Same as (a) but for the numerically exact solutions. (c) Same as (a) but for the SU(3) FP-TWA results. The number of the generated Langevin trajectories is 200 000200000200\;000200 000.
@@
When the hopping amplitude is zero, i.e., $J=0$, the SU(3) dTWA reproduces the exact quantum time evolution generated by the master equation with a direct product state for the typical initial state. This is attributed to the linearity of the classical equation of motion in the interaction $U$ term and the dissipation $\gamma$ term as implied in Eq. (10). Figure 8(a) displays numerical results calculated by the SU(3) dTWA for a single-site system, demonstrating that this specific semiclassical representation successfully reproduces the exact solutions of the dynamics of $\langle{\hat{X}}_{3}(t)\rangle={\rm Tr}[\rho_{\rm eff}(t)T_{3}]$ for arbitrary dissipation strength at all time.
However, as shown in Fig. 8(b), the quantitative validity of the SU(3) FP-TWA is verified only within a finite timescale. This limitation is due to the nonlinearity in the dissipation term of the Langevin equation. For $\gamma=0.5$, the SU(3) FP-TWA result coincides well with the numerically exact solution until $tU/\hbar\sim 3$. After this timescale, the SU(3) FP-TWA result deviates from the numerically exact solution and finally converges to a false steady state with $\langle{\hat{X}}_{3}(t)\rangle<1$ for $tU/\hbar\gg 1$. Note that the exact steady state for the present setup is the trivial bosonic vacuum state $|0\rangle$, implying $\langle{\hat{X}}_{3}(t)\rangle=1$ for $tU/\hbar\gg 1$. For larger values of $\gamma$, i.e., $\gamma=1$, $2$, and $4$, the threshold time is estimated as $tU/\hbar\sim 1.5$, $1$, and $0.5$, respectively. Hence, the valid timescale of this approximation shortens with increasing $\gamma$. Notice also that the time-evolved states for these different values of $\gamma$ result in the same false steady state at long time. Comparing the results in Figs. 8(a) and  8(b), the local linearity of the dissipative force in the SU(3) dTWA is essential to reproduce the correct steady state reached via the successive onsite losses over time.
@@
Reflecting these consequences, we also find in Fig. 10(a) that the two-body loss rate $\kappa$ extracted from the dTWA results exhibits a maximum around at $\gamma\sim 2$, although the values themselves are smaller than the exact ones. However, the agreement with the three-dimensional experiment in Fig. 2(c) suggests that the deviations between these two results of the loss rate may be smaller in higher dimensions because the valid timescale of the dTWA increases with the connectivity between sites, i.e., the coordination number $z=2d$ Kunimi et al. (2021). Additionally, we perform the power-law fitting for the long-time behavior of the density in Fig. 9. Figures 10(c)–10(e) also demonstrate that the atomic density $n_{j}(t)$, obtained from three different methods, fit well with the power-law function in the long-time regime, clearly distinguishing it from the short-time behavior, which is well fitted by $f_{\kappa}(t)=(1+\kappa t)^{-1}$. This qualitative difference is also evidence when compared to the mean-field results shown in the same figures. Notably, as shown in Fig. 10(b), the rescaled power exponent $c/c_{\rm max}$ in the power-law function obtained from the SU(3) dTWA closely follow the exact results for varying $\gamma$, while the FP-TWA results deviate significantly.
The right panels of Fig. 9 show the SU(3) FP-TWA results of the atomic density $n_{j}(t)$ compared with the numerically exact solutions. For the weak dissipation strength, i.e., $\gamma=0.1$, the SU(3) FP-TWA result deviates form the exact solution around $tU/\hbar=2$. However, for the strong dissipation strength such as $\gamma=4$ and $5$, the SU(3) FP-TWA and exact results coincide rather well until $tU/\hbar\lesssim 6$. We should note that this coincidence is not controlled in this approximation. Although the SU(3) FP-TWA can capture the continuous Zeno effect for the strong dissipation strength, it estimates the threshold at $\gamma\sim 1$, which is smaller than the exact solution. As shown in Fig. 10(a), the maximum of the two-body loss rate $\kappa$ appears around the same value of $\gamma$. Comparing the SU(3) dTWA and FP-TWA results, we conclude that the qualitative performance of the SU(3) FP-TWA is worse than that of the SU(3) dTWA at long time.
@@ V Conclusions
In conclusions, we have generalized the SU(${\cal N}$) TWA method for the Bose-Hubbard system to include the Lindblad dynamics in the presence of onsite dissipation terms. We introduced two semiclassical representations of the dynamics governed by the effective Lindblad master equation: the discrete TWA representation and the Fokker-Planck TWA representation. These methods were applied to the dissipative Bose-Hubbard system with the two-body losses, previously studied experimentally in Ref. Tomita et al. (2017). We demonstrated that these semiclassical representations can qualitatively capture the essential features of the continuous quantum Zeno effect in the dissipative nonunitary dynamics, as observed experimentally. In particular, the discrete TWA representation was found to produce reasonable thresholds for the dissipation strength, distinguishing between the Zeno and non-Zeno regimes, which are comparable to the experimental results of the analog quantum simulation Tomita et al. (2017). This is attributed to the onsite linearity of the classical equation of motion in the discrete TWA representation, which was used in the Monte Carlo simulations. The local linearity is highly powerful because a collection of fluctuating trajectories with locally linear equations results in a power-law decay in the long-time atomic density loss dynamics, a feature that cannot be captured by the naive mean-field approximation.
The general strategy that we have developed opens a new avenue for addressing a wide range of open quantum many-body systems, particularly in two and three dimensions, which are computationally challenging. As a future study, it would be interesting to apply those SU(${\cal N}$) TWA methods to other experiments associated with dissipative Bose-Hubbard systems Labouvie et al. (2016); Bouganne et al. (2020) and discuss the reproducibility of nontrivial driven-dissipative dynamical phenomena reported therein. The generalization of the methods to SU(${\cal N}$) spin systems for ${\cal N}>3$ is straightforward, and these methods potentially enable one to evaluate the impact of dissipation on various experiments of analog quantum simulation, such as optical tweezer arrays loaded with Rydberg gases, trapped ions, and large-$S$ dipolar gases in optical lattices. Beyond the present capability of the methods based on onsite variables, it is also a promising direction to incorporate the ideas of the cluster TWA with fluctuating cluster variables Wurtz et al. (2018) to treat offsite quantum correlations caused by the hopping term more accurately Nagao and Yunoki (2024).
@@ Acknowledgements.
Parts of the numerical simulations were carried out on the HOKUSAI supercomputing system at RIKEN (Project ID No. Q22576) and the supercomputer Fugaku installed at the RIKEN Center for Computational Science. We thank Kazuhiro Seki for his helpful comments on MPI parallelization for Monte Carlo simulations. We also thank Masaya Kunimi and Tetsuro Nikuni for fruitful discussions. We acknowledge Takafumi Tomita and Yoshiro Takahashi for sharing the experimental data in Ref. Tomita et al. (2017). This work was financially supported by JSPS KAKENHI (Grants No. JP18H05228, No. JP21H01014, No. JP21H04446, and No. JP21H03455), by MEXT Q-LEAP (Grant No. JPMXS0118069021), by JST FOREST (Grant No. JPMJFR202T), and by JST COI-NEXT (Grant No. JPMJPF2221). This work was also supported by the RIKEN TRIP initiative (RIKEN Quantum).